# Extended-Use Designs on Very Large Online Platforms

### v1

@@ 4.4. Lulling Users
Autopopulated Fields. Interfaces also lull users into continued engagement by autopopulating text fields and offering one-click interaction mechanisms. Feeding the user pre-populated text both models platform-desired behavior for the user and makes it possible for the user to enact this behavior with a single tap or keystroke. For example, LinkedIn proactively suggests possible job-search queries based on the user’s profile, clicking Amazon’s search bar brings up a list of suggested shopping searches, and TikTok suggests the user respond to a post by offering several pre-selected emojis to choose from (Figure 12).
Figure 12. Autopopulated responses and searches lull the user into taking action with almost no thought required. Left: TikTok serves the user possible responses to another person. Right: Amazon (top) and LinkedIn (bottom) fill out search suggestions.
@@
5. Results Part 2: Prevalence and Distribution of EUDs on VLOPs
After distilling and organizing the patterns described above, we systematically coded all 17 VLOP websites and their 16 corresponding apps for the presence of these designs. In doing so, we both documented the specific features they used to invoke these EUD patterns and the frequency of each type of pattern across the ecosystem of VLOPs. Further supplementary material, including interactive visualizations of this data, a database of feature instances, and our codebook is available here.
5.1. The Prevalence of EUDs across VLOPs
We found that EUDs were extremely common on VLOPs, and 100% of VLOPs contained features that employed one or more of the EUD patterns described above (see Figure 15). Across the 16 VLOP app interfaces, we identified 334 instances of features that employ EUD patterns, and across the 17 VLOP website interfaces, we identified 249 instances of features that employ EUD patterns. However, the use of EUDs was not evenly distributed, and for example, we found that TikTok’s app interface contains 36 different features that draw on EUD patterns, while the Apple App Store app uses only nine. On average, VLOP apps ($N=16$) used $mean=20.94$ EUD features ($SD=8.10$, $min=9$, $max=36$, $median=20$), with TikTok, Instagram, Snapchat, Facebook, and LinkedIn each using more than 25. The prevalance of EUDs was similar on VLOP websites ($N=17$), where platforms employed an average of $mean=14.65$ features ($SD=6.75$, $min=3$, $max=26$, $median=13$). Website version of LinkedIn, Tiktok, YouTube, Instagram, and X topped the list, each including more than 20 features that employ EUDs.
Figure 15. The Prevalence of EUDs by Platforms; app(left), website(right)
@@ 5.1. The Prevalence of EUDs across VLOPs
Figure 16. The Prevalence of EUDs by Platform Type(eCommerce, information, social media); app(left), website(right)
5.2. The Prevalence of each Type of EUD
@@
Figure 17. Average Number of EUD Features per platform; app(left), website(right)
Figure 18. EUD Pattern Occurrences on VLOPs; app(left), website(right)
We also looked at the prevalence of each of these four strategies across the three types of VLOPs (social media, eCommerce, information). We found that social media platforms (apps: $mean=19.00,sd=7.55$; websites: $mean=8.38,sd=2.13$) generally use pressuring users and lulling users (apps: $mean=29.00,sd=14.01$; websites: $mean=7.63,sd=0.92$) as their main strategies. In contrast, e-commerce platforms primarily use pressuring users (apps: $mean=13.00,sd=7.55$; websites: $mean=6.17,sd=1.94$) and lulling users (apps: $mean=16.00,sd=14.01$; websites: $mean=4.83,sd=2.23$). Information platforms tend to use trapping users (apps: $mean=2.00,sd=7.21$; websites: $mean=3.00,sd=1.41$) and pressuring users (apps: $mean=4.00,sd=7.55$; websites: $mean=5.50,sd=2.12$), though they generally use fewer engagement strategies overall compared to social media and e-commerce platforms. Figure 19 shows the tree diagram of each VLOP’s usage of EUDs.
Figure 19. Tree diagrams of the EUDs on each VLOP app. Each node represents a different EUD, and each color represents a different strategy (yellow: pressuring users; green: enticing users; red: trapping users; blue: lulling users). As shown here, eCommerce platforms and information platforms are visibly sparser than social media platforms, which use twice as many EUDs. Interactive versions of these visualizations that allow for exploring the EUDs used by each app can be found here
6. Results Part 3: Interface Vignettes Combining EUDs
Here, we complement our data describing EUDs and their prevalence with illustrations of their use in context. We provide three different vignettes that characterize the holistic experience of encountering EUDs as a VLOP user and show how multiple EUDs can work together, each compounding the effects of the others.
@@ 6.1. Vignette 1: Creating Chain Reactions
We observed that VLOPs often funnel users through multiple screens that each incorporate one or more EUDs. These chain reactions create a behavioral funnel that guides users step-by-step through a series of interactions, with each encouraging prolonged engagement by nudging them into the step that follows.
For example (see storyboard in Figure 20), LinkedIn creates an unread notification badge as a visual cue to trigger Obligatory Engagement. This creates a subtle sense of urgency or discomfort, prompting the user to click on the badge to clear it. Upon clicking the badge, the user encounters a prompt encouraging them to congratulate a friend on achieving a milestone, creating a subtle Artificial Social Obligation without overt pressure. When the user clicks on the prompt encouraging them to “Say congrats,” they are taken to a chat interface with an Autopopulated Field supplying the text to send, requiring only one click to send a congratulatory message. Sending this message then triggers EUDs for the recipient: the recipient receives a social activity alert in the form of an in-app badge, push notification, or email, informing them of the message. This, in turn, creates another incomplete task that requires action to click the button, and generate social pressure to respond.
With this set of EUD features, the platform creates a complete feedback loop: even though both features—a notification badge and prompts to congratulate a friend—rely on mechanisms that subtly pressure the user, their combination also creates a nuanced experience of lulling the users into a state where they are more likely to follow a pre-designed path of actions. This gently guides them into extended engagement while subduing the feeling of being pressured or manipulated to stay on the platform. Low-friction designs like autopopulated congratulatory text further lull the user into extended engagement. And this chain reaction moves across the network of connected individuals, with one user’s actions triggering EUDs in another user’s interface.
Figure 20. Creating Chain Reactions: LinkedIn creates chain reaction by having one EUD feature lead to another.
@@ 6.2. Vignette 2: Bombarding Users via Multiple Lines of Attack
In other cases, VLOPs employ multiple EUDs separately and simultaneously, bombarding the user with many parallel attempts to hijack their attention. For example, on YouTube’s main page, features based on lulling mechanisms, like infinite scroll and autoplay, provide a continuous flow of content, keeping users passively engaged without requiring active effort. Features based on pressuring mechanisms, such as unread badges, subtly push users to resolve notifications at the same time. Features based on enticing mechanisms, like provocative thumbnails, instantly grab attention, even as the user scrolls through options or clears notifications. Features based on trapping mechanisms, like picture-in-picture playback, keep users connected to content even as they maintain other threads of engagement. These overlapping features work together to cumulatively create an environment that aggressively pursue the user’s attention.
Figure 21. Bombarding the user: YouTube displays a range of EUDs representing all four strategies on the main page.
6.3. Vignette 3: Fusing EUDs
Finally, we saw that VLOPs often merge two or more EUD patterns into a single UI element to enhance its potential for extending user engagement. For example, Booking.com presents the user with a dialog encouraging them to complete a series of bookings with the platform within a certain time frame (see Figure 22). In doing so, they create a Platform-Manufactured Task, giving the user a to-do list of bookings, and they exert Fabricated Time Pressure, giving the user a time constraint which adds urgency to their engagement. The user’s progress is represented by five circles, which break down the broader goal (completing five bookings) into smaller, manageable tasks (one booking at a time). By merging these two EUDs, a single UI element can leverage multiple approaches at once to increasing the user’s engagement.
Figure 22. Fusing EUDs: Booking.com creates a Platform-Manufactured Task with Fabricated Time Pressure, employing multiple forms of pressure through a single feature.
@@ 7. Discussion
7.1. The Ubiquity and Aggressiveness of EUDs
@@
We found that EUDs are ubiquitous across VLOPs, which serve hundreds of millions of users worldwide. These design categories have been invented by platforms to encourage prolonged use and re-engagement, and thus contribute to displacement of other pursuits and lead users to blame themselves for excessive time spent with engaging interfaces online. We found that platforms: 1) pressure users into engaging by creating artificial social obligations and task lists, 2) entice users by leveraging reward- and pleasure-mediated mechanisms that often involve curiosity teasing and tailoring, 3) trap users through disorientating interface elements that cause confusion, block navigation, and inhibit decision-making, and 4) lull users via automated, low-friction designs that reduce user self-determination by automating decision-making, overloading the UI with heuristic-driven stimuli, and encouraging mindlessness.
@@
These categories of manipulation work together to both guide users toward decisions that extend their use and away from decisions that end their use. Enticement designs attempt to make users feel excited to stay engaged while pressureful designs make them feel bad about leaving. Trapping designs make leaving difficult, while lulling designs make staying easy. VLOPs leverage these strategies in combination, linking them together to form chain reactions, bombarding the user with a multitude of competing EUDs, and fusing multiple EUDs into single features. Together, these forces create a current that carries the user along in one direction, steering them away from disengagement decisions while simultaneously funneling them into engagement ones.
7.2. A Conceptual Model of Extended-Use Design
@@
This taxonomy builds upon several pre-existing taxonomies that have described dark patterns, attention capture, and designs that promote compulsive use (Flayelle et al., 2023; Monge Roffarello et al., 2023; Radesky et al., 2022) and situates them on VLOPs, the most commonly used platforms with the greatest global reach. Its novel contributions to this academic field include its combination of theory-driven and empirically driven data to create a taxonomy that describes the motives behind design, rather than only describing designs phenomenologically. In other words, while dark pattern typologies describing specific interface designs such as “confirmshaming” may come and go, we aimed to describe the underlying purpose and psychological mechanisms by which users pay attention to digital products, use them for extended periods of time, return after non-use, or feel compelled to re-engage.
@@
Figure 23. A conceptual model of the common strategies underlying EUDs. Some EUDs manipulate emotions (e.g., manufacturing feelings of guilt, anxiety, or stress by pressuring users; manufacturing feelings of curiosity, pleasure, or popularity by enticing users). Some EUDs manipulate cognitive load (e.g., making navigation difficult by trapping users; making navigation simple by lulling users). These designs work to reward users emotionally and cognitively for continuing to engage and punishing users emotionally and cognitively for attempting to disengage.
@@
Our findings show that EUDs operate on two levels (see Figure 23). First, they leverage users’ emotional responses, manufacturing guilt, anxiety, curiosity, and pleasure in ways that encourage engagement and discourage disengagement. Second, they manipulate cognitive load, making engagement decisions frictionless and disengagement decisions frictionful. In both cases, EUDs offer rewards for behaviors that extend engagement and punish behaviors that end it. This model offers some potential insight into future regulatory strategies; for example, punishing designs that invoke negative emotional experiences and increase cognitive load may be better suited to regulation than those that reward users for extended engagement. The parallel structure that we repeatedly encountered—wherein engagement flows (e.g., signing up for a service) lulled users into continuing, and disengagement flows (e.g., cancelling a service) trapped users such that they struggled to make forward progress—also suggest a need for regulation. Requiring that engagement and disengagement experiences have equal levels of lulling and trapping would support users in making more autonomous decisions about their time and attention.
@@ 7.3. The Challenge of “Gray Patterns”
However, prior work documents that it is deeply problematic to assume that users’ engagement behaviors always reflect their preferences (Richards and Hartzog, 2024). Using extended use as a measure of users’ interest in a product is confounded by the very fact that EUDs intend to manufacture extended use itself. And it is clear that attention-economy incentives drive the proliferation of EUDs: in our sample, social media platforms that depend on user attention employed twice as many EUDs as platforms that rely on other business models. This suggests both a need for regulation and landscape that defies it, leaving regulators with the challenge of curtailing overly aggressive or invasive EUDs while leaving space for products to engage their users in ways they find pleasurable and meaningful.
Historically, dark patterns have been identified by their explicit deception. In the case of EUDs, users may be deceived primarily by their own lack of awareness of their susceptibility to these patterns. Our results show that EUDs are pervasive and, if effective, have the potential to dramatically reshape how people spend their time. This suggests a need for regulations that go beyond straightforward cases of deception and put bounds on companies’ ability to subtly manipulate users via gray patterns (that are concerning but not quite “dark”).

### v2

@@ 4.4. Lulling Users
Autopopulated Fields. Interfaces also lull users into continued engagement by autopopulating text fields and offering one-click interaction mechanisms. Feeding the user pre-populated text both models platform-desired behavior for the user and makes it possible for the user to enact this behavior with a single tap or keystroke. For example, Instagram proactively suggests possible entertainment searches based on the user’s profile, clicking Amazon’s search bar brings up a list of suggested shopping searches, and TikTok suggests the user respond to a post by offering several pre-selected emojis to choose from (Figure 12).
Figure 12. Autopopulated responses and searches lull the user into taking action with almost no thought required. Left: TikTok serves the user possible responses to another person. Right: Amazon (top) and Instagram (bottom) fill out search suggestions.
@@
5. Results Part 2: Prevalence and Distribution of EPDs on VLOPs
After distilling and organizing the patterns described above, we systematically coded all 17 VLOP websites and their 16 corresponding apps for the presence of these designs. In doing so, we both documented the specific features they used to invoke these EPD patterns and the frequency of each type of pattern across the ecosystem of VLOPs. Further supplementary material, including interactive visualizations of this data, a database of feature instances, and our codebook is available on the Engagement-Prolonging Design website [^5] https://engagementprolongingdesign.github.io/EngagementProlongingDesign/.
5.1. The Prevalence of EPDs across VLOPs
We found that EPDs were extremely common on VLOPs, and 100% of VLOPs contained features that employed one or more of the EPD patterns described above (see Figure 15). Across the 16 VLOP app interfaces, we identified 334 instances of features that employ EPD patterns, and across the 17 VLOP website interfaces, we identified 249 instances of features that employ EPD patterns. However, the use of EPDs was not evenly distributed, and for example, we found that TikTok’s app interface contains 36 different features that draw on EPD patterns, while the Apple App Store app uses only nine. On average, VLOP apps ($N=16$) used $mean=20.94$ EPD features ($SD=8.10$, $min=9$, $max=36$, $median=20$), with TikTok, Instagram, Snapchat, Facebook, and LinkedIn each using more than 25. The prevalance of EPDs was similar on VLOP websites ($N=17$), where platforms employed an average of $mean=14.65$ features ($SD=6.75$, $min=3$, $max=26$, $median=13$). Website version of LinkedIn, Tiktok, YouTube, Instagram, and X topped the list, each including more than 20 features that employ EPDs.
Figure 15. The Prevalence of EPDs by Platforms; app(left), website(right)
@@ 5.1. The Prevalence of EPDs across VLOPs
Figure 16. The Prevalence of EPDs by Platform Type(eCommerce, information, social media); app(left), website(right)
5.2. The Prevalence of each Type of EPD
@@
Figure 17. Average Number of EPD Features per platform; app(left), website(right)
Figure 18. EPD Pattern Occurrences on VLOPs; app(left), website(right)
We also looked at the prevalence of each of these four strategies across the three types of VLOPs (social media, eCommerce, information). We found that social media platforms (apps: $mean=19.00,sd=7.55$; websites: $mean=8.38,sd=2.13$) generally use pressuring users and lulling users (apps: $mean=29.00,sd=14.01$; websites: $mean=7.63,sd=0.92$) as their main strategies. In contrast, e-commerce platforms primarily use pressuring users (apps: $mean=13.00,sd=7.55$; websites: $mean=6.17,sd=1.94$) and lulling users (apps: $mean=16.00,sd=14.01$; websites: $mean=4.83,sd=2.23$). Information platforms tend to use trapping users (apps: $mean=2.00,sd=7.21$; websites: $mean=3.00,sd=1.41$) and pressuring users (apps: $mean=4.00,sd=7.55$; websites: $mean=5.50,sd=2.12$), though they generally use fewer engagement strategies overall compared to social media and e-commerce platforms. Figure 7 shows the tree diagram of each VLOP’s usage of EPDs.
Figure 19. Tree diagrams of the EPDs on each VLOP app. Each node represents a different EPD, and each color represents a different strategy (yellow: pressuring users; green: enticing users; red: trapping users; blue: lulling users). As shown here, eCommerce platforms and information platforms are visibly sparser than social media platforms, which use twice as many EPDs. Interactive versions of these visualizations that allow for exploring the EPDs used by each app can be found on the Engagement-Prolonging Design website [^7] https://engagementprolongingdesign.github.io/EngagementProlongingDesign/.
6. Results Part 3: Interface Vignettes Combining EPDs
Here, we complement our data describing EPDs and their prevalence with illustrations of their use in context. We provide three different vignettes that characterize the holistic experience of encountering EPDs as a VLOP user and show how multiple EPDs can work together, each compounding the effects of the others.
@@ 6.1. Vignette 1: Creating Chain Reactions
We observed that VLOPs often funnel users through multiple screens that each incorporate one or more EPDs. These chain reactions create a behavioral funnel that guides users step-by-step through a series of interactions, with each encouraging prolonged engagement by nudging them into the step that follows.
For example (see storyboard in Figure 20), LinkedIn creates an unread notification badge as a visual cue to trigger Obligatory Engagement. This creates a subtle sense of urgency or discomfort, prompting the user to click on the badge to clear it. Upon clicking the badge, the user encounters a prompt encouraging them to congratulate a friend on achieving a milestone, creating a subtle Artificial Social Obligation without overt pressure. When the user clicks on the prompt encouraging them to “Say congrats,” they are taken to a chat interface with an Autopopulated Field supplying the text to send, requiring only one click to send a congratulatory message. Sending this message then triggers EPDs for the recipient: the recipient receives a social activity alert in the form of an in-app badge, push notification, or email, informing them of the message. This, in turn, creates another incomplete task that requires action to click the button, and generate social pressure to respond.
With this set of EPD features, the platform creates a complete feedback loop: even though both features—a notification badge and prompts to congratulate a friend—rely on mechanisms that subtly pressure the user, their combination also creates a nuanced experience of lulling the users into a state where they are more likely to follow a pre-designed path of actions. This gently guides them into extended engagement while subduing the feeling of being pressured or manipulated to stay on the platform. Low-friction designs like autopopulated congratulatory text further lull the user into extended engagement. And this chain reaction moves across the network of connected individuals, with one user’s actions triggering EPDs in another user’s interface.
Figure 20. Creating Chain Reactions: LinkedIn creates chain reaction by having one EPD feature lead to another.
@@ 6.2. Vignette 2: Bombarding Users via Multiple Lines of Attack
In other cases, VLOPs employ multiple EPDs separately and simultaneously, bombarding the user with many parallel attempts to hijack their attention. For example, on YouTube’s main page, features based on lulling mechanisms, like infinite scroll and autoplay, provide a continuous flow of content, keeping users passively engaged without requiring active effort. Features based on pressuring mechanisms, such as unread badges, subtly push users to resolve notifications at the same time. Features based on enticing mechanisms, like provocative thumbnails, instantly grab attention, even as the user scrolls through options or clears notifications. Features based on trapping mechanisms, like picture-in-picture playback, keep users connected to content even as they maintain other threads of engagement. These overlapping features work together to cumulatively create an environment that aggressively pursue the user’s attention.
Figure 21. Bombarding the user: YouTube displays a range of EPDs representing all four strategies on the main page.
6.3. Vignette 3: Fusing EPDs
Finally, we saw that VLOPs often merge two or more EPD patterns into a single UI element to enhance its potential for extending user engagement. For example, Booking.com presents the user with a dialog encouraging them to complete a series of bookings with the platform within a certain time frame (see Figure 22). In doing so, they create a Platform-Manufactured Task, giving the user a to-do list of bookings, and they exert Fabricated Time Pressure, giving the user a time constraint which adds urgency to their engagement. The user’s progress is represented by five circles, which break down the broader goal (completing five bookings) into smaller, manageable tasks (one booking at a time). By merging these two EPDs, a single UI element can leverage multiple approaches at once to increasing the user’s engagement.
Figure 22. Fusing EPDs: Booking.com creates a Platform-Manufactured Task with Fabricated Time Pressure, employing multiple forms of pressure through a single feature.
@@ 7. Discussion
7.1. The Ubiquity and Aggressiveness of EPDs
@@
We found that EPDs are ubiquitous across VLOPs, which serve hundreds of millions of users worldwide. These design categories have been invented by platforms to encourage prolonged use and re-engagement, and thus contribute to displacement of other pursuits and lead users to blame themselves for excessive time spent with engaging interfaces online. We found that platforms: 1) pressure users into engaging by creating artificial social obligations and task lists, 2) entice users by leveraging reward- and pleasure-mediated mechanisms that often involve curiosity teasing and tailoring, 3) trap users through disorientating interface elements that cause confusion, block navigation, and inhibit decision-making, and 4) lull users via automated, low-friction designs that reduce user self-determination by automating decision-making, overloading the UI with heuristic-driven stimuli, and encouraging mindlessness.
@@
These categories of manipulation work together to both guide users toward decisions that extend their use and away from decisions that end their use. Enticement designs attempt to make users feel excited to stay engaged while pressureful designs make them feel bad about leaving. Trapping designs make leaving difficult, while lulling designs make staying easy. VLOPs leverage these strategies in combination, linking them together to form chain reactions, bombarding the user with a multitude of competing EPDs, and fusing multiple EPDs into single features. Together, these forces create a current that carries the user along in one direction, steering them away from disengagement decisions while simultaneously funneling them into engagement ones.
7.2. A Conceptual Model of Engagement-Prolonging Design
@@
This taxonomy builds upon several pre-existing taxonomies that have described dark patterns, attention capture, and designs that promote compulsive use (Flayelle et al., 2023; Monge Roffarello et al., 2023; Radesky et al., 2022; Gray et al., 2024) and situates them on VLOPs, the most commonly used platforms with the greatest global reach. Its novel contributions to this academic field include its combination of theory-driven and empirically driven data to create a taxonomy that describes the motives behind design, rather than only describing designs phenomenologically. In other words, while dark pattern typologies describing specific interface designs such as “confirmshaming” may come and go, we aimed to describe the underlying purpose and psychological mechanisms by which users pay attention to digital products, use them for extended periods of time, return after non-use, or feel compelled to re-engage.
@@
Figure 23. A conceptual model of the common strategies underlying EPDs. Some EPDs manipulate emotions (e.g., manufacturing feelings of guilt, anxiety, or stress by pressuring users; manufacturing feelings of curiosity, pleasure, or popularity by enticing users). Some EPDs manipulate cognitive load (e.g., making navigation difficult by trapping users; making navigation simple by lulling users). These designs work to reward users emotionally and cognitively for continuing to engage and punishing users emotionally and cognitively for attempting to disengage.
@@
Our findings show that EPDs operate on two levels (see Figure 23). First, they leverage users’ emotional responses, manufacturing guilt, anxiety, curiosity, and pleasure in ways that encourage engagement and discourage disengagement. Second, they manipulate cognitive load, making engagement decisions frictionless and disengagement decisions frictionful. In both cases, EPDs offer rewards for behaviors that extend engagement and punish behaviors that end it. This model offers some potential insight into future regulatory strategies; for example, punishing designs that invoke negative emotional experiences and increase cognitive load may be better suited to regulation than those that reward users for extended engagement. The parallel structure that we repeatedly encountered—wherein engagement flows (e.g., signing up for a service) lulled users into continuing, and disengagement flows (e.g., canceling a service) trapped users such that they struggled to make forward progress—also suggests a need for regulation. Requiring that engagement and disengagement experiences have equal levels of lulling and trapping would support users in making more autonomous decisions about their time and attention.
@@ 7.3. The Challenge of “Gray Patterns”
However, prior work documents that it is deeply problematic to assume that users’ engagement behaviors always reflect their preferences (Richards and Hartzog, 2024). Using extended use as a measure of users’ interest in a product is confounded by the very fact that EPDs intend to manufacture extended use itself. And it is clear that attention-economy incentives drive the proliferation of EPDs: in our sample, social media platforms that depend on user attention employed twice as many EPDs as platforms that rely on other business models. This suggests both a need for regulation and landscape that defies it, leaving regulators with the challenge of curtailing overly aggressive or invasive EPDs while leaving space for products to engage their users in ways they find pleasurable and meaningful.
Historically, dark patterns have been identified by their explicit deception, but engagement has avoided regulation because of the slipperiness of its definition between metric, ideology, and harmful practice (Richards and Hartzog, 2024). In the case of EPDs, users may be deceived primarily by their own lack of awareness of their susceptibility to these patterns (Luguri and Strahilevitz, 2021; M. Bhoot et al., 2020). Our results show that EPDs are pervasive and, if effective, have the potential to dramatically reshape how people spend their time. This suggests a need for regulations that go beyond straightforward cases of deception and put bounds on companies’ ability to subtly manipulate users via gray patterns (that are concerning but not quite “dark”).